OpenStudy (tylermckinney16):

Sammy is a sales clerk in a bicycle shop. He's paid $6.25/hour for a 37.5 hour week, plus a commission of 6% of his sales for the week. In one week Sammy's sales were $2319.75. Calculate Sammy's total earnings for the week. Show work.

OpenStudy (danjs):

His salary is money from hourly work plus 6% of the sales he did. He worked at $6.25 an hour for 37.5 hours, that totals 6.25*37.5 = $234.38 6% of the total sales 0.06 * 2319.75 = $139.19 So the total he earned is 234.38 + 139.19 = $373.57